OpenStudy (anonymous):

Which sequences have a common ratio of 1/3? Choose exactly two answers that are correct.

OpenStudy (anonymous):

the one where the second term is one third of the first

OpenStudy (anonymous):

I believe it's the 2nd and 3rd ones. :) @satellite73 \[-3, 9 ,-1,\frac{ 1 }{ 3 },-\frac{ 1 }{ 9 },...\] \[-\frac{ 1 }{ 2 },-\frac{ 1 }{ 6 },-\frac{ 1 }{ 18 },-\frac{ 1 }{ 54 },-\frac{ 1 }{ 162 },...\] \[900,300,100,\frac{ 100 }{ 3 },\frac{ 100 }{ 9 },...\] \[100,300,900,2700,8100\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

second one for sure

OpenStudy (anonymous):

also the third